# Well quasi-order and atomicity for combinatorial structures under consecutive orders

### v3

@@ 7. Well quasi-order for bountiful structures
Let $C=\operatorname{Av}(B)$ be an avoidance set of a poset of bountiful structures. The following are equivalent:
@@
In each of the following posets, an avoidance set is wqo if and only if it is finite:
@@ 8. Atomicity for Bountiful Structures
Let $C=\operatorname{Av}(B)$ be an avoidance set of a poset of valid structures under the consecutive order. Then $C$ is atomic if and only if:
@@
Suppose that $\Gamma_{C}$ is a bicycle but is not strongly connected and, aiming for a contradiction, suppose that it has an ambiguous path $\pi$. We can extend $\pi$ to an ambiguous path $\pi^{\prime}$ which enters more than one strongly connected component. Take two distinct structures $\sigma_{1},\sigma_{2}\in\pi^{\prime}$. Since $C$ is atomic, there exists $\theta\in C$ such that $\sigma_{1},\sigma_{2}\leq\theta$, so $\Pi(\sigma_{1}),\Pi(\sigma_{2})\leq\Pi(\theta)$ by Proposition 4.3. Since $\sigma_{1},\sigma_{2}$ are distinct, they cannot embed in $\theta$ via the same embedding. Therefore, they correspond to different subsets of the points of $\theta$, so their paths are different subpaths of $\Pi(\theta)$. In other words, $\Pi(\theta)$ has two distinct subpaths which are both isomorphic to $\pi^{\prime}$. This means that $\Pi(\theta)$ traverses every edge of $\pi^{\prime}$ twice, contradicting the assumption that $\pi^{\prime}$ enters more than one strongly connected component. Hence $\Gamma_{C}$ cannot have ambiguous paths. ∎
@@
Let $C=\operatorname{Av}(B)$ be an avoidance set of a poset of bountiful structures under the consecutive order. Then $C$ is atomic if and only if:
@@
$(\Rightarrow)$ Since bountiful posets are valid, if $C$ is atomic then by Theorem 8.1, $(2)$ holds and $\Gamma_{B}$ is strongly connected or a bicycle with no ambiguous paths. Since $C$ is bountiful, all paths in $\Gamma_{B}$ are ambiguous, so $\Gamma_{B}$ must be strongly connected, giving $(1)$.
@@
Consider the conditions of Corollary 8.2. It is easy construct $\Gamma_{C}$ and check whether it is strongly connected. For the second condition, there are only finitely many elements in $C_{[1,b-1]}$ and finitely many elements in $C_{b}$ that they could be contained in. So both conditions are decidable, giving the result. ∎
@@
In of each of the following posets, an avoidance set $C$ is atomic if and only if its factor graph is strongly connected and for each $\sigma\in C_{[1,b-1]}$ there is $\rho\in C_{b}$ such that $\sigma\leq\rho$.
@@
It is easy to see that condition $(2)$ of Theorem 8.1 is decidable. For condition $(1)$, it is decidable whether the factor graph is strongly connected or a bicycle. If it is a bicycle, by assumption, it is decidable whether it contains ambiguous paths. Hence, it is decidable whether $(X,\leq)$ is atomic. ∎
@@ 9. Words under the consecutive order
An avoidance set $C$ of $\mathcal{W}_{A}$ is atomic if and only if the following hold:
@@ 10. An invalid example: double ascents
In this section we will introduce an example of invalid strcutures: permutations with at most one inversion between consecutive entries. These will be referred to as double ascents.
@@
A consecutive inversion in a permutation $\sigma=\sigma_{1}\dots\sigma_{n}$ is a pair $(\sigma_{i},\sigma_{i+1})$ of consecutive entries such that $\sigma_{i+1}\leq_{v}\sigma_{i}$.
@@
A permutation $\sigma=\sigma_{1}\dots\sigma_{n}$ is a double ascent if it contains at most one consecutive inversion. It is a single ascent if it contains no consecutive inversions.
We will refer to increasing permutations as ascents, so single and double ascents are permutations containing at most 1 and 2 ascents as consecutive subpermutations respectively.
@@
$213\cong 324\leq 135246$ under the consecutive order. Here the underlying contiguous mapping with respect to $\leq_{v}$ is $f:[1,3]\rightarrow[1,6]$ defined by $f(x)=x+1$.
@@
Consider the double ascents 13245 and 12453. We see that $13245\hskip-0.85358pt\!\upharpoonright_{[3,5]}=345\cong 123=12453\hskip-0.85358pt\!\upharpoonright_{[1,3]}$. However, if we identify these points, we obtain the permutation 1324675, which has two consecutive inversions: those given by the subsequences 32 and 75. Hence 1324675 is not a double ascent. ∎
@@ 11. Double ascents and words
To begin, we will describe the relationship between double ascents and words over $\{r,l\}$. The idea here is that we assign entries of double ascents to the right (‘$r$’) or left (‘$l$’) according to whether they come to the right or left of the consecutive inversion.
@@
Let $w$ be a word over $\{l,r\}$. The associated permutation of $w$ is the double ascent $\textbf{A}(w)$ formed from the empty permutation as follows. Reading left to right through $w$, for each letter we add a point to the permutation, in increasing value order. If we read the letter $l$, we add the point to the rightmost position on the left of the second point of the consecutive inversion, and if we read the letter $r$ we add the point to the rightmost position on the right of the first point of the consecutive inversion.
In this way, the points to the left of the second point of the consecutive inversion (those assigned ‘left’) are the points in the first ascent. Similarly, points to the right of the first point of the consecutive inversion (those assigned ‘right’) are the points of the second ascent.
@@
We will construct $\textbf{A}(llrlr)$. For clarity, we will draw a bar through the position of the consecutive inversion. The first two letters are $l$, so the first two entries of $\textbf{A}(llrlr)$ by value are $12|$. The next entry is given by the letter $r$, so this will be added to the right of the consecutive inversion, giving $12|3$. Following this, the next entry is to the left of the consecutive inversion, giving $124|3$. Finally, the last letter $r$ gives the largest entry by value to the right of the consecutive inversion, so we obtain $124|35$. Removing the bar, $\textbf{A}(llrlr)=12435$.
@@
By inverting the process described in Definition 11.1, for any double ascent $\sigma$ we can find the words in $\textbf{W}(\sigma)$. To do this, begin with the empty word. We will then consider the entries $\sigma$ in order of their values, smallest to largest, and for each entry add a letter to construct words in $\textbf{W}(\sigma)$. If an entry of $\sigma$ is to the left of the second entry of the consecutive inversion, we will add the letter $l$. On the other hand, if an entry is to the right of the first entry of the consecutive inversion, we will add the letter $r$. It can be seen that the resulting words are in $\textbf{W}(\sigma)$ by constructing $\textbf{A}(w)$ for each word $w$ and observing that this is precisely $\sigma$.
@@
$(\Leftarrow)$ Let $\sigma$ be a single ascent of length $n\geq 1$. Then the words $l^{n},r^{n}\in\textbf{W}(\sigma)$, so $|\textbf{W}(\sigma)|>1$. $(\Rightarrow)$ We prove the contrapositive. Let $\sigma$ be a double ascent, but not a single ascent. Then the process to find the words in $\textbf{W}(\sigma)$ yields exactly one word, as each letter is determined uniquely by its position relative to the consecutive inversion in $\sigma$. ∎
@@
$(\Rightarrow)$ Consider constructing words in $\textbf{W}(\sigma)$. Initially, we may start with either the letter $l$ or $r$. Once we have added the letter $r$, every following letter must also be $r$, since the consecutive subword $rl$ would correspond to a consecutive inversion. Hence, the possible words are precisely those in the set $\{l^{a}r^{c}\hskip 2.84526pt|\hskip 2.84526pta+c=n\}$. $(\Leftarrow)$ It can easily be checked that the associated permutation of any word in $\{l^{a}r^{c}\hskip 2.84526pt|\hskip 2.84526pta+c=n\}$ is $\sigma$ (the single ascent of length $n$). ∎
@@
Since $\sigma\leq\rho$, it follows that $\sigma\cong\rho\hskip-0.85358pt\!\upharpoonright_{[k,\hskip 1.99168ptk+|\sigma|-1]}$ for some $k\in[1,|\rho|]$. Since neither $\sigma$ nor $\rho$ is a single ascent, $\textbf{W}(\sigma)$ and $\textbf{W}(\rho)$ are uniquely defnied, so $\textbf{W}(\sigma)=\textbf{W}(\rho\hskip-0.85358pt\!\upharpoonright_{[k,\hskip 1.99168ptk+|\sigma|-1]})\leq\textbf{W}(\rho)$. ∎
@@ 12. WQO for double ascents
An avoidance set $C=\operatorname{Av}(B)$ of $\mathcal{DA}$ is wqo if and only if the avoidance set $K=\operatorname{Av}(\textbf{W}(B))$ of $\mathcal{W}_{\{l,r\}}$ is wqo, i.e. $\Gamma_{K}$ contains no in-out cycles.
@@ 13. Atomicity for double ascents
If an avoidance set $C=\operatorname{Av}(B)$ of double ascents contains non-extendable single ascents, the following are equivalent:
@@
If an avoidance set $C=\operatorname{Av}(B)$ of double ascents contains no non-extendable single ascents, then $C$ is atomic if and only if the following hold:
@@
If follows from (i) and Proposition 3.13 that $\Gamma_{K}$ is atomic. Therefore there exists a path $\pi$ in $\Gamma_{K}$ such that $\Pi(\textbf{W}(\sigma)),\Pi(\textbf{W}(\rho))\leq\pi$. By Proposition 9.5, $\textbf{W}(\sigma),\textbf{W}(\rho)\leq\Sigma(\pi)$ in $K$. We observe that $\sigma,\rho\leq\textbf{A}(\Sigma(\pi))$ in $C$ by Lemma 11.10. Therefore $C$ satisfies the JEP and so is atomic.
@@
Now we will prove that (i) holds. Again, suppose to the contrary that $\Gamma_{K}$ is neither strongly connected nor a bicycle. By Proposition 3.13, the set of paths of $\Gamma_{K}$ is not atomic, so we may take paths $\eta,\pi$ in $\Gamma_{K}$ which do not join. It follows that $\Sigma(\eta),\Sigma(\pi)$ do not join in $K$ by Proposition 9.5. By Lemma 12.3, $\textbf{A}(\Sigma(\eta)),\textbf{A}(\Sigma(\pi))\in C$, so by assumption we may extend them to double ascents $A_{\eta}^{+},A_{\pi}^{+}\in C$ respectively which are not single ascents. Since $C$ is assumed to be atomic, there exists $\theta\in C$ such that $A_{\eta}^{+},A_{\pi}^{+}\leq\theta$. We know that $\textbf{W}(\theta)\in K$ by Lemma 12.3 and $\textbf{W}(A_{\eta}^{+}),\textbf{W}(A_{\pi}^{+})\leq\textbf{W}(\theta)$ by Lemma 11.8. Since $\Sigma(\eta)\leq\textbf{W}(A_{\eta}^{+})$ and $\Sigma(\pi)\leq\textbf{W}(A_{\pi}^{+})$, this implies that $\Sigma(\eta),\Sigma(\pi)\leq\textbf{W}(\theta)$, yielding a contradiction as $\Sigma(\eta)$ and $\Sigma(\pi)$ do not join. We conclude that $\Gamma_{K}$ must be strongly connected or a bicycle. ∎
@@
Let $C=\operatorname{Av}(B)$ be a finitely based avoidance set of double ascents under the consecutive order and $K=\operatorname{Av}(\textbf{W}(B))$. Then $C$ is atomic if and only if the following hold:
@@
The conditions of Theorem 13.9 can be tested algorithmically, so the result follows. ∎

### v4

@@ 7. Well quasi-order for bountiful structures
Let $C=\operatorname{Av}(B)$ be a finitely based avoidance set of a poset of bountiful structures. The following are equivalent:
@@
If $\Gamma_{C}$ is empty, the result is trivial, so we assume that this is not the case for the remainder of this proof.
@@
In each of the following posets, a finitely based avoidance set is wqo if and only if it is finite:
@@ 8. Atomicity for Bountiful Structures
Let $C=\operatorname{Av}(B)$ be an infinite, finitely based avoidance set of a poset of valid structures under the consecutive order. Then $C$ is atomic if and only if:
@@
Suppose that $\Gamma_{C}$ is a bicycle but is not strongly connected and, aiming for a contradiction, suppose that it has an ambiguous path $\pi$. By adding new vertices to the start or end of $\pi$, we can extend $\pi$ to an ambiguous path $\pi^{\prime}$ which enters more than one strongly connected component. Take two distinct structures $\sigma_{1},\sigma_{2}\in\pi^{\prime}$. Since $C$ is atomic, there exists $\theta\in C$ such that $\sigma_{1},\sigma_{2}\leq\theta$, so $\Pi(\sigma_{1}),\Pi(\sigma_{2})\leq\Pi(\theta)$ by Proposition 4.3. Since $\sigma_{1},\sigma_{2}$ are distinct, they cannot embed in $\theta$ via the same embedding. Therefore, they correspond to different subsets of the points of $\theta$, so their paths are different subpaths of $\Pi(\theta)$. In other words, $\Pi(\theta)$ has two distinct subpaths which are both isomorphic to $\pi^{\prime}$. This means that $\Pi(\theta)$ traverses every edge of $\pi^{\prime}$ twice, contradicting the assumption that $\pi^{\prime}$ enters more than one strongly connected component. Hence $\Gamma_{C}$ cannot have ambiguous paths. ∎
@@
Let $C=\operatorname{Av}(B)$ be an infinite, finitely based avoidance set of a poset of bountiful structures under the consecutive order. Then $C$ is atomic if and only if:
@@
$(\Rightarrow)$ Since bountiful posets are valid, if $C$ is atomic then by Theorem 8.1, $(2)$ holds and $\Gamma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C}}$ is strongly connected or a bicycle with no ambiguous paths. Since $C$ is bountiful, all paths in $\Gamma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C}}$ are ambiguous, so $\Gamma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C}}$ must be strongly connected, giving $(1)$.
@@
It is easy to check whether a finitely based avoidance set is finite, as we can just check whether its factor graph contains cycles. For the finite avoidance sets, we can algorithmically check for atomicity. For infinite avoidance sets, consider the conditions of Corollary 8.2. It is easy construct $\Gamma_{C}$ and check whether it is strongly connected. For the second condition, there are only finitely many elements in $C_{[1,b-1]}$ and finitely many elements in $C_{b}$ that they could be contained in. So both conditions are decidable, giving the result. ∎
@@
In of each of the following posets, an infinite, finitely based avoidance set $C$ is atomic if and only if its factor graph is strongly connected and for each $\sigma\in C_{[1,b-1]}$ there is $\rho\in C_{b}$ such that $\sigma\leq\rho$.
@@
As in the proof of Corollary 8.2, it is decidable whether an avoidance set is finite, and whether a finite avoidance set is atomic. For the infinite avoidance sets, it is easy to see that condition $(2)$ of Theorem 8.1 is decidable. For condition $(1)$, it is decidable whether the factor graph is strongly connected or a bicycle. If it is a bicycle, by assumption, it is decidable whether it contains ambiguous paths. Hence, it is decidable whether $(X,\leq)$ is atomic. ∎
@@ 9. Words under the consecutive order
An infinite avoidance set $C$ of $\mathcal{W}_{A}$ is atomic if and only if the following hold:
@@ 10. An invalid example: double ascents
In this section we will introduce an example of invalid structures: permutations with at most one inversion between consecutive entries. These will be referred to as double ascents.
@@
An adjacent inversion in a permutation $\sigma=\sigma_{1}\dots\sigma_{n}$ is a pair $(\sigma_{i},\sigma_{i+1})$ of consecutive entries such that $\sigma_{i+1}\leq_{v}\sigma_{i}$.
@@
A permutation $\sigma=\sigma_{1}\dots\sigma_{n}$ is a double ascent if it contains at most one adjacent inversion. It is a single ascent if it contains no adjacent inversions.
We will refer to increasing permutations as ascents, so single and double ascents are permutations containing no ascents and at most 1 ascent as consecutive subpermutations respectively.
@@
$213\cong 324\leq 135246$ under the consecutive order (which is defined with respect to the value linear order, rather than the more usual position order, as discussed above). Here the underlying contiguous mapping with respect to $\leq_{v}$ is $f:[1,3]\rightarrow[1,6]$ defined by $f(x)=x+1$.
@@
Consider the double ascents 13245 and 12453. We see that $13245\hskip-0.85358pt\!\upharpoonright_{[3,5]}=345\cong 123=12453\hskip-0.85358pt\!\upharpoonright_{[1,3]}$. However, if we identify these points, we obtain the permutation 1324675, which has two adjacent inversions: those given by the subsequences 32 and 75. Hence 1324675 is not a double ascent. ∎
@@ 11. Double ascents and words
To begin, we will describe the relationship between double ascents and words over $\{r,l\}$. The idea here is that we assign entries of double ascents to the right (‘$r$’) or left (‘$l$’) according to whether they come to the right or left of the adjacent inversion.
@@
Let $w$ be a word over $\{l,r\}$. The associated permutation of $w$ is the double ascent $\textbf{A}(w)$ formed from the empty permutation as follows. Reading left to right through $w$, for each letter we add a point to the permutation, in increasing value order. If we read the letter $l$, we add the point to the rightmost position on the left of the second point (with respect to $\leq_{p}$) of the adjacent inversion, and if we read the letter $r$ we add the point to the rightmost position on the right of the first point (again with respect to $\leq_{p}$) of the adjacent inversion.
In this way, the points to the left of the second point of the adjacent inversion (those assigned ‘left’) are the points in the first ascent. Similarly, points to the right of the first point of the adjacent inversion (those assigned ‘right’) are the points of the second ascent.
@@
We will construct $\textbf{A}(llrlr)$. For clarity, we will draw a bar through the position of the adjacent inversion. The first two letters are $l$, so the first two entries of $\textbf{A}(llrlr)$ by value are $12|$. The next entry is given by the letter $r$, so this will be added to the right of the adjacent inversion, giving $12|3$. Following this, the next entry is to the left of the adjacent inversion, giving $124|3$. Finally, the last letter $r$ gives the largest entry by value to the right of the adjacent inversion, so we obtain $124|35$. Removing the bar, $\textbf{A}(llrlr)=12435$.
@@
By inverting the process described in Definition 11.1, for any double ascent $\sigma$ we can find the words in $\textbf{W}(\sigma)$. To do this, begin with the empty word. We will then consider the entries $\sigma$ in order of their values, smallest to largest, and for each entry add a letter to construct words in $\textbf{W}(\sigma)$. If an entry of $\sigma$ is to the left of the second entry of the adjacent inversion, we will add the letter $l$. On the other hand, if an entry is to the right of the first entry of the adjacent inversion, we will add the letter $r$. It can be seen that the resulting words are in $\textbf{W}(\sigma)$ by constructing $\textbf{A}(w)$ for each word $w$ and observing that this is precisely $\sigma$.
@@
$(\Leftarrow)$ Let $\sigma$ be a single ascent of length $n\geq 1$. Then the words $l^{n},r^{n}\in\textbf{W}(\sigma)$, so $|\textbf{W}(\sigma)|>1$. $(\Rightarrow)$ We prove the contrapositive. Let $\sigma$ be a double ascent, but not a single ascent. Then the process to find the words in $\textbf{W}(\sigma)$ yields exactly one word, as each letter is determined uniquely by its position relative to the adjacent inversion in $\sigma$. ∎
@@
$(\Rightarrow)$ Consider constructing words in $\textbf{W}(\sigma)$. Initially, we may start with either the letter $l$ or $r$. Once we have added the letter $r$, every following letter must also be $r$, since the consecutive subword $rl$ would correspond to an adjacent inversion. Hence, the possible words are precisely those in the set $\{l^{a}r^{c}\hskip 2.84526pt|\hskip 2.84526pta+c=n\}$. $(\Leftarrow)$ It can easily be checked that the associated permutation of any word in $\{l^{a}r^{c}\hskip 2.84526pt|\hskip 2.84526pta+c=n\}$ is $\sigma$ (the single ascent of length $n$). ∎
@@
Since $\sigma\leq\rho$, it follows that $\sigma\cong\rho\hskip-0.85358pt\!\upharpoonright_{[k,\hskip 1.99168ptk+|\sigma|-1]}$ for some $k\in[1,|\rho|]$. Since neither $\sigma$ nor $\rho$ is a single ascent, $\textbf{W}(\sigma)$ and $\textbf{W}(\rho)$ are uniquely defined, so $\textbf{W}(\sigma)=\textbf{W}(\rho\hskip-0.85358pt\!\upharpoonright_{[k,\hskip 1.99168ptk+|\sigma|-1]})\leq\textbf{W}(\rho)$. ∎
@@ 12. WQO for double ascents
A finitely based avoidance set $C=\operatorname{Av}(B)$ of $\mathcal{DA}$ is wqo if and only if the avoidance set $K=\operatorname{Av}(\textbf{W}(B))$ of $\mathcal{W}_{\{l,r\}}$ is wqo, i.e. $\Gamma_{K}$ contains no in-out cycles.
@@ 13. Atomicity for double ascents
If an infinite avoidance set $C=\operatorname{Av}(B)$ of double ascents contains non-extendable single ascents, the following are equivalent:
@@
If an infinite avoidance set $C=\operatorname{Av}(B)$ of double ascents contains no non-extendable single ascents, then $C$ is atomic if and only if the following hold:
@@
It follows from (i) and Proposition 3.13 that the set of paths of $\Gamma_{K}$ is atomic. Therefore there exists a path $\pi$ in $\Gamma_{K}$ such that $\Pi(\textbf{W}(\sigma)),\Pi(\textbf{W}(\rho))\leq\pi$. By Proposition 9.5, $\textbf{W}(\sigma),\textbf{W}(\rho)\leq\Sigma(\pi)$ in $K$. We observe that $\sigma,\rho\leq\textbf{A}(\Sigma(\pi))$ in $C$ by Lemma 11.10. Therefore $C$ satisfies the JEP and so is atomic.
@@
Now we will prove that (i) holds. Again, suppose to the contrary that $\Gamma_{K}$ is neither strongly connected nor a bicycle. By Proposition 3.13, the set of paths of $\Gamma_{K}$ is not atomic, so we may take paths $\eta,\pi$ in $\Gamma_{K}$ which do not join. It follows that $\Sigma(\eta),\Sigma(\pi)$ do not join in $K$ by Proposition 9.5. Therefore, $\Sigma(\eta),\Sigma(\pi)$ cannot both be associated with single ascents, because their paths would be in the same left-right bicycle, and so would join. If needed, extend $\Sigma(\eta)$ to be the word of a double ascent in $C$ which is not a single ascent, call this $w$. Note that $\Sigma(\eta)\leq w$. Since $C$ is assumed to be atomic, there exists $\theta\in C$ such that $\textbf{A}(w),\textbf{A}(\Sigma(\pi))\leq\theta$. Therefore, by Lemma 11.8, $w,\Sigma(\pi)\leq\textbf{W}(\theta)$. Hence $\Sigma(\eta),\Sigma(\pi)\leq\textbf{W}(\theta)$, yielding a contradiction as $\Sigma(\eta)$ and $\Sigma(\pi)$ do not join. We conclude that $\Gamma_{K}$ must be strongly connected or a bicycle.
@@
Let $C=\operatorname{Av}(B)$ be an infinite, finitely based avoidance set of double ascents under the consecutive order and $K=\operatorname{Av}(\textbf{W}(B))$. Then $C$ is atomic if and only if the following hold:
@@
We can determine whether $C$ is finite by checking whether the factor graph $\Gamma_{K}$ of $K=\operatorname{Av}(\textbf{W}(B))$ contains cycles. If $C$ is finite, atomicity can be determined algorithmically. If $C$ is infinite, the conditions of Theorem 13.9 can be tested algorithmically, so the result follows. ∎